# Benchmarking Retrieval-Augmented Generation in Multi-Modal Contexts

### v1

@@ Abstract
This paper introduces Multi-Modal Retrieval-Augmented Generation (M2RAG), a benchmark designed to evaluate the effectiveness of Multi-modal Large Language Models (MLLMs) in leveraging knowledge from multi-modal retrieval documents. The benchmark comprises four tasks: image captioning, multi-modal question answering, multi-modal fact verification, and image reranking. All tasks are set in an open-domain setting, requiring RAG models to retrieve query-relevant information from a multi-modal document collection and use it as input context for RAG modeling. To enhance the context utilization capabilities of MLLMs, we also introduce Multi-Modal Retrieval-Augmented Instruction Tuning (MM-RAIT), an instruction tuning method that optimizes MLLMs within multi-modal contexts. Our experiments show that MM-RAIT improves the performance of RAG systems by enabling them to effectively learn from multi-modal contexts. All data and code are available at https://github.com/NEUIR/M2RAG.
Benchmarking Retrieval-Augmented Generation in Multi-Modal Contexts
Zhenghao Liu1, Xingsheng Zhu1, Tianshuo Zhou1, Xinyi Zhang1, Xiaoyuan Yi2, Yukun Yan3, Yu Gu1, Ge Yu1, Maosong Sun3thanks:   indicates corresponding author. 1Department of Computer Science and Technology, Northeastern University, China 2Microsoft Research Asia, Beijing, China 3Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China
1 Introduction
With the rapid development of Large Language Models (LLMs), such as GPT-4 OpenAI (2023) and LLaMA Touvron et al. (2023), they have demonstrated strong emergent abilities in many NLP tasks Wei et al. (2022); Zhao et al. (2023). However, LLMs often face the issue of hallucinations, making them produce unreliable responses Ji et al. (2023); Huang et al. (2023); Shuster et al. (2021). Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Asai et al. (2024b); Shi et al. (2024); Yao et al. (2023) has proven effective in mitigating this hallucination problem by integrating external knowledge with LLMs.
@@ 1 Introduction
Figure 1: Illustration of Multi-Modal RAG Tasks. We incorporate multi-modal retrieval documents as the input context for MLLMs.
@@
To enhance LLMs with retrieved knowledge, existing approaches typically feed retrieved documents as input contexts, prompting LLMs to generate responses based on this in-context information Ram et al. (2023). Existing RAG approaches Petroni et al. (2021); Lin et al. (2024) usually focus on retrieving textual knowledge from corpora to aid LLMs in answering queries. Recent studies Hu et al. (2024); Sharifymoghaddam et al. (2024) have extended RAG to Multi-modal Large Language Models (MLLMs), enabling them to address knowledge-intensive and information-seeking tasks that involve visual queries. However, these approaches largely rely on text or images as the sole sources of external knowledge, often overlooking the critical role of multi-modal data in providing richer and more comprehensive information that leads to producing more accurate answers Hu et al. (2024); Liu et al. (2023b).
@@
To advance RAG modeling in multi-modal scenarios, we introduce the Multi-Modal RAG (M2RAG) benchmark, designed to explore the effectiveness of MLLMs by feeding multi-modal retrieved documents as the input contexts to answer the question. As shown in Figure 1, we can use images or text as queries to retrieve multi-modal documents via multi-modal dense retrievers Liu et al. (2023b); Zhou et al. (2024b, a). These multi-modal documents are then used as the input contexts to assist MLLMs during generation. Specifically, our M2RAG benchmark includes four distinct tasks: image captioning, multi-modal question answering, multi-modal fact verification, and image reranking. Different from existing works Aghajanyan et al. (2022); Sharifymoghaddam et al. (2024), M2RAG is built upon high-quality datasets Chang et al. (2022); Mishra et al. (2022) and designs four evaluation tasks in the open-domain setting, aiming to assess the effectiveness of MLLMs in leveraging the knowledge from multi-modal contexts.
@@
In this paper, we also propose the Multi-Modal Retrieval Augmented Instruction Tuning (MM-RAIT) method to adapt MLLMs to the multi-modal in-context learning scenario, enhancing the effectiveness of MLLMs in utilizing the knowledge from these multi-modal retrieval documents. Specifically, we design task-specific prompt templates and fine-tune MLLMs on the M2RAG benchmark, making MLLMs maintain contextual awareness during generation. Our experimental results demonstrate that using retrieved knowledge significantly enhances MLLMs’ performance, achieving significant improvements in both zero-shot and few-shot settings. After training with MM-RAIT, MiniCPM-V and Qwen2-VL show an average improvement of 27% and 34% over vanilla RAG modeling methods, showing the effectiveness of MM-RAIT.
2 Related Work
@@
Existing RAG models Shi et al. (2024); Asai et al. (2024a); Yu et al. (2023b); Yan et al. (2024) typically rely on dense retrievers Karpukhin et al. (2020); Xiong et al. (2021a); Ren et al. (2021); Xiong et al. (2021b); Gao and Callan (2022) or sparse retrievers like BM25 Robertson et al. (2009) for text document retrieval. More recent efforts have integrated multi-modal retrieval methods, allowing the inclusion of rich external knowledge from different modalities within RAG frameworks. For example, some works Liu et al. (2023b); Zhou et al. (2024b) have introduced unified multi-modal retrieval systems that map images and texts into a shared semantic space. This approach allows for single-modal matching, cross-modal matching, and modality routing within the embedding space. VISTA Zhou et al. (2024a) further enhances multi-modal retrieval by optimizing synthetic training data and refining training strategies. These advancements enable the retrieval of multi-modal knowledge, providing a way for evaluating the effectiveness of MLLMs in multi-modal contexts.
@@
Multi-modal Large Language Models (MLLMs) Achiam et al. (2023); Team et al. (2023); Sun et al. (2024b, a); Aghajanyan et al. (2022); Lu et al. (2024) have proven their effectiveness in understanding, integrating, and utilizing both visual and textual knowledge in generation tasks. Models like BLIP Li et al. (2022, 2023), LLaVA Liu et al. (2023a), and Flamingo Alayrac et al. (2022) build the MLLMs by combining pre-trained vision encoders with Large Language Models (LLMs) to process multi-modal inputs for generation. Thriving on the advancements in MLLMs, researchers pay more attention to extending the advantages of Retrieval-Augmented Generation (RAG) to these MLLMs, enhancing their generation capability.
@@
Figure 2: Examples of Different Tasks Defined in the M2RAG Benchmark. All tasks are designed for the open-domain setting. Thus, we present the input, ground truth answers, and retrieved documents for each task.
Multi-modal RAG has demonstrated its potential to enhance knowledge-intensive and information-seeking tasks, such as question answering Chang et al. (2022); Marino et al. (2019) and fact verification Mishra et al. (2022). These models utilize retrieval-based multi-modal documents to provide richer and contextually relevant information. Additionally, other works have applied multi-modal RAG to improve the performance of MLLMs on the tasks like image captioning Lin et al. (2014); Young et al. (2014) and generation Yasunaga et al. (2023); Yu et al. (2023a); Sharifymoghaddam et al. (2024). However, existing multi-modal benchmarks Johnson et al. (2017); Schuhmann et al. (2021); Lin et al. (2014); Young et al. (2014); Marino et al. (2019) are typically tailored to specific tasks and lack a comprehensive framework for evaluating multi-modal RAG systems.
3 M2RAG Benchmark for Multi-Modal Retrieval-Augmented Generation
@@
In this section, we describe our Multi-Modal Retrieval-Augmented Generation (M2RAG) benchmark. We first introduce the RAG tasks in M2RAG, followed by a detailed explanation of the construction process. Finally, we present a comparison of existing multi-modal benchmarks with M2RAG.
@@
Task Definition. As shown in Figure 2, M2RAG defines four tasks to evaluate the capabilities of MLLMs in open-domain RAG scenarios: image captioning, multi-modal question answering, multi-modal fact verification, and image reranking. For each task, MLLMs are required to retrieve knowledge from the multi-modal document collection $\mathcal{D}$ and generate responses to answer the question $q$. Details of the prompt templates of different tasks are shown in Appendix A.3.
@@
Image Captioning Task. Image Captioning is a widely used task for evaluating the performance of multi-modal RAG models Aghajanyan et al. (2022); Sharifymoghaddam et al. (2024). In this task, an image is provided as the query $q$, and the document collection $\mathcal{D}$ is constructed using image documents that contain captions. The goal of image captioning is to generate concise and semantically coherent captions that accurately describe the image content. Unlike previous works Aghajanyan et al. (2022); Sharifymoghaddam et al. (2024), we source image captions from WebQA Chang et al. (2022), where all image documents are collected from Wikimedia Commons. These captions often include important details, such as named entities, which make the task more challenging and provide crucial information for query matching Liu et al. (2023b). More comparison details of the image captioning tasks in different benchmarks are shown in Appendix A.4.
Benchmarks Input Modality Knowledge Source Retrieval Modality Multi Task Open Domain
MSCOCO Lin et al. (2014) Image Web [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Flickr30K Young et al. (2014) Image Web [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
K-VQA Shah et al. (2019) Multi Wikipedia [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WebQA Chang et al. (2022) Text Wikipedia Multi [Uncaptioned image] [Uncaptioned image]
MRAG-Bench Hu et al. (2024) Multi Web and Other Sources Image [Uncaptioned image] [Uncaptioned image]
M2RAG (Ours) Multi Wikipedia, Twitter Multi [Uncaptioned image] [Uncaptioned image]
Table 1: Comparison of Multi-Modal Benchmarks.
@@
Multi-Modal Question Answering Task. Following the WebQA benchmark Chang et al. (2022), the Multi-Modal QA task involves answering text-based queries $q$ by leveraging both text and image documents. The document collection $\mathcal{D}$ consists of both text and image documents containing captions. Additionally, we extend WebQA to an open-domain setting by instructing the retriever to return query-relevant documents from the collection $\mathcal{D}$, as demonstrated by Liu et al. (2023b).
@@
Multi-Modal Fact Verification Task. The Multi-Modal Fact Verification task challenges MLLMs to verify the accuracy of claims using multi-modal evidence. In this task, the query $q$ can be a multi-modal claim, and the document collection $\mathcal{D}$ consists of both text and image documents, where the image documents do not contain captions. The relationship between the claim and the evidence is categorized into three possible outcomes: “Support”, “Refute”, or “Insufficient”, indicating whether the evidence supports, refutes, or lacks sufficient information to verify the claim. We build this task on the Factify dataset Mishra et al. (2022), but we focus on open-domain fact verification by retrieving evidence from a multi-modal document collection Thorne et al. (2018).
@@
Image Reranking Task. In the Image Reranking task, the objective is to identify the most relevant images based on a given image description. Here, the image description serves as the query $q$, and the document collection $\mathcal{D}$ consists of image documents that do not include captions. For each description, we first use a multi-modal retriever to retrieve image candidates based solely on their image features and then rerank the images using MLLMs. To adapt MLLMs for this task, we follow prior work Muennighoff (2022) and compute the Perplexity (PPL) score for reranking image candidates based on their image features. This approach models the relevance between queries and images in a manner similar to image captioning, where a lower PPL score indicates greater relevance between the candidate image and the given query.

### v2

@@ Abstract.
With the rapid advancement of Multi-modal Large Language Models (MLLMs), their capability in understanding both images and text has greatly improved. However, their potential for leveraging multi-modal contextual information in Retrieval-Augmented Generation (RAG) remains largely underexplored. To address this gap, this paper introduces Multi-Modal Retrieval-Augmented Generation (M2RAG), a benchmark designed to evaluate the effectiveness of Multi-modal Large Language Models in leveraging knowledge from multi-modal retrieval documents. The benchmark comprises four tasks: image captioning, multi-modal question answering, multi-modal fact verification, and image reranking. All tasks are set in an open-domain setting, requiring RAG models to retrieve query-relevant information from a multi-modal document collection and use it as contextual input for RAG modeling. To enhance the context utilization capabilities of MLLMs, we also introduce Multi-Modal Retrieval-Augmented Instruction Tuning (MM-RAIT), an instruction tuning method that optimizes MLLMs within multi-modal contexts. Our experiments demonstrate the effectiveness of MM-RAIT by significantly improving the quality of responses generated by different RAG models, outperforming MiniCPM-V 2.6 and Qwen2-VL with 34% and 33% gains, respectively. All data and code are available at https://github.com/NEUIR/M2RAG.
submissionid: 123-A56-BU3submissionid: 4874
1. Introduction
@@
Figure 1. Illustration of Multi-Modal RAG Tasks. The documents, retrieved from different multi-modal knowledge sources, are used as contextual input to MLLMs.
With the rapid development of Large Language Models (LLMs), such as GPT-4 (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have demonstrated strong emergent abilities in many natural language processing (NLP) tasks (Wei et al., 2022; Zhao et al., 2023). However, LLMs often face the issue of hallucinations, causing them to produce unreliable responses (Ji et al., 2023; Huang et al., 2023; Shuster et al., 2021). Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Asai et al., 2024b; Shi et al., 2024; Yao et al., 2023) has proven effective in mitigating this hallucination problem by incorporating external knowledge into the generation process, thereby improving the factual accuracy and reliability of LLM outputs.
@@
To enhance LLMs with retrieved knowledge, existing approaches typically feed retrieved documents into LLMs as input contexts, prompting them to generate responses based on this in-context information (Ram et al., 2023). Existing RAG approaches (Petroni et al., 2021; Lin et al., 2024) usually focus on retrieving textual knowledge from corpora to aid LLMs in answering queries. Recent studies (Hu et al., 2024; Sharifymoghaddam et al., 2024; Chen et al., 2022; Caffagni et al., 2024; Ding et al., 2024; Cui et al., 2024) have extended Retrieval-Augmented Generation (RAG) to Multi-modal Large Language Models (MLLMs), enabling them to handle knowledge-intensive and information-seeking tasks involving visual queries. However, most existing benchmarks route queries to different models in an oracle manner, and evaluate MLLMs using either text or images as the sole external knowledge source. In contrast, real-world RAG scenarios often require retrieving query-relevant information from sources of diverse modalities (Liu et al., 2023b), and effectively integrating complementary signals across modalities to generate accurate answers.
@@
To advance RAG modeling in multi-modal scenarios, we introduce the Multi-Modal RAG (M2RAG) benchmark, designed to explore the effectiveness of MLLMs by feeding multi-modal retrieved documents as the input contexts to answer the question. As shown in Figure 1, we can first use images or text as queries to retrieve multi-modal documents via multi-modal dense retrievers (Liu et al., 2023b; Zhou et al., 2024b, a). Then these multi-modal documents are used as the input contexts to assist MLLMs to generate responses for the user query. Our benchmark emphasizes the integration of visual and textual modalities during both retrieval and generation stages. Different from existing works (Aghajanyan et al., 2022; Sharifymoghaddam et al., 2024), M2RAG is constructed based on high-quality datasets (Chang et al., 2022; Mishra et al., 2022) and introduces four diverse tasks, image captioning, multi-modal question answering, multi-modal fact verification, and image reranking, for evaluation. To better reflect real-world scenarios, these tasks are reformulated in an open-domain setting, enabling a more comprehensive assessment of how effectively MLLMs can leverage knowledge from multi-modal contexts.
@@
In this paper, we also propose the Multi-Modal Retrieval Aug-mented Instruction Tuning (MM-RAIT) method to adapt MLLMs to the multi-modal in-context learning scenario, enhancing the effectiveness of MLLMs in utilizing the knowledge from these multi-modal retrieval documents. Specifically, we design task-specific prompt templates for different tasks in the M2RAG benchmark and then fine-tune MLLMs within multi-modal retrieved context, making MLLMs maintain contextual awareness during generation. Our experimental results demonstrate that using retrieved knowledge significantly enhances MLLMs’ performance, achieving significant improvements in both zero-shot and few-shot settings. After training with MM-RAIT, MiniCPM-V and Qwen2-VL show an average improvement of 34% and 33% over vanilla RAG modeling methods, showing the effectiveness of MM-RAIT.
2. Related Work
@@
Existing RAG models (Shi et al., 2024; Asai et al., 2024a; Yu et al., 2023b; Yan et al., 2024) typically rely on dense retrievers (Karpukhin et al., 2020; Xiong et al., 2021b; Ren et al., 2021; Xiong et al., 2021a; Gao and Callan, 2022) or sparse retrievers (Robertson et al., 2009) for text document retrieval. Recent works (Liu et al., 2023b; Zhou et al., 2024b, a) usually focus on broadening the effectiveness of text retriever to multi-modal retrieval scenarios, allowing the inclusion of rich external knowledge from different modalities within RAG frameworks. They build unified multi-modal retrieval systems that map images and texts into a shared semantic space, which allows for single-modal matching, cross-modal matching, and modality routing (Liu et al., 2023b). These advancements enable the retrieval of multi-modal knowledge, providing a way for evaluating the effectiveness of MLLMs within multi-modal contexts.
@@
Multi-modal Large Language Models (MLLMs) (Achiam et al., 2023; Team et al., 2023; Sun et al., 2024b, a; Aghajanyan et al., 2022; Lu et al., 2024) have proven their effectiveness in understanding, integrating, and utilizing both visual and textual knowledge in generation tasks. Models like BLIP (Li et al., 2022, 2023), LLaVA (Liu et al., 2023a), and Flamingo (Alayrac et al., 2022) build MLLMs by combining pre-trained vision encoders with Large Language Models (LLMs), enabling LLMs to process multi-modal inputs during generation. Emu2 (Sun et al., 2024a) further extends the generative potential of MLLMs by pretraining using a large-scale multi-modal corpus with a unified autoregressive objective, thereby improving the model’s transferability to a wide range of downstream tasks. Qwen-VL (Wang et al., 2024) further enhances multi-modal understanding and generation by integrating a high-resolution visual encoder with a fine-grained, multi-stage fusion mechanism, effectively aligning visual tokens with linguistic representations. In parallel, MiniCPM-V (Yao et al., 2024) adopts a lightweight vision encoder and a compact LLM, offering a favorable trade-off between performance and efficiency, particularly suited for deployment in resource-constrained environments. Thriving on the advancements in MLLMs, researchers pay more attention to extending the advantages of Retrieval-Augmented Generation (RAG) to these MLLMs, enhancing their generation capability using the knowledge from different modalities.
@@
Figure 2. Examples of Different Tasks Defined in the M2RAG Benchmark. All tasks are designed for the open-domain setting. Thus, we present the input, ground truth answers, and retrieved documents for each task.
Multi-modal RAG has demonstrated its potential to enhance knowledge-intensive and information-seeking tasks, such as question answering (Chang et al., 2022; Marino et al., 2019) and fact verification (Mishra et al., 2022). These works (Chen et al., 2022; Caffagni et al., 2024; Ding et al., 2024; Yu et al., 2024) utilize retrieval-based multi-modal documents to provide richer and contextually relevant information. Wiki-LLaVA (Caffagni et al., 2024) enhances performance on Visual Question Answering (VQA) tasks by retrieving external knowledge based on the input image. VisRAG (Yu et al., 2024) further extends to document-level VQA tasks by directly leveraging document page images instead of extracted text, preserving and utilizing the original data within documents to enhance MLLM generation. RA-BLIP (Ding et al., 2024) introduces an Adaptive Selective Knowledge Generation (ASKG) strategy, which enables the generator to autonomously assess the relevance of retrieved knowledge, thereby achieving strong denoising performance and effectively reducing the interference of irrelevant information during retrieval and generation. Additionally, several studies have applied multi-modal RAG to improve the performance of MLLMs on tasks like image captioning (Lin et al., 2014; Young et al., 2014; Hu et al., 2023) and generation (Yasunaga et al., 2023; Yu et al., 2023a; Sharifymoghaddam et al., 2024). MORE (Cui et al., 2024) further extends multi-modal RAG to commonsense reasoning tasks. During training, it introduces a “query dropout” method to prevent the language model from either completely ignoring the retrieved multi-modal documents due to potential noise or becoming overly reliant on possibly noisy retrieval results. While recent works have made significant progress in applying multi-modal RAG to various tasks, the evaluation of these systems remains underexplored. Existing multi-modal benchmarks (Johnson et al., 2017; Schuhmann et al., 2021; Lin et al., 2014; Young et al., 2014; Marino et al., 2019) are typically tailored to specific tasks and lack a unified framework for systematically assessing the performance of multi-modal RAG models.
3. M2RAG Benchmark for Multi-Modal Retrieval-Augmented Generation
@@
In this section, we introduce our Multi-Modal Retrieval-Augmented Generation (M2RAG) benchmark. We first introduce the RAG tasks included in M2RAG, followed by a detailed explanation of the construction process. Finally, we provide a comparative analysis between existing multi-modal benchmarks and M2RAG.
@@
Task Definition. As shown in Figure 2, M2RAG defines four tasks to evaluate the capabilities of MLLMs in open-domain RAG scenarios: image captioning, multi-modal question answering, multi-modal fact verification, and image reranking. For each task, MLLMs are required to retrieve knowledge from the multi-modal document collection $\mathcal{D}$ and generate responses to answer the question $q$.
@@
Image Captioning Task. Image Captioning is a widely used task for evaluating the performance of multi-modal RAG models (Aghajanyan et al., 2022; Sharifymoghaddam et al., 2024). In this task, an image is provided as the query $q$, and the document collection $\mathcal{D}$ is constructed using image documents that contain captions. The goal of image captioning is to generate concise and semantically coherent captions that accurately describe the image content. Unlike previous works (Aghajanyan et al., 2022; Sharifymoghaddam et al., 2024), we collect image captions from WebQA (Chang et al., 2022), where all image documents are collected from Wikimedia Commons. These captions often include important details for verbalizing the semantics of images, such as named entities, making the task more challenging (Liu et al., 2023b).
@@
Multi-Modal Question Answering Task. Multi-Modal Question Answering (QA) is a task for assessing the capabilities of multi-modal RAG models in understanding and reasoning across both textual and visual modalities (Chang et al., 2022). Given a textual query $q$, the model aims to generate accurate and informative answers by retrieving and leveraging relevant documents from a multi-modal collection $\mathcal{D}$, which includes text and image documents with captions. We follow WebQA benchmark (Chang et al., 2022) and extend it to an open-domain setting, where the retriever selects query-relevant documents from the entire collection $\mathcal{D}$, following Liu et al. (2023b).
@@
Multi-Modal Fact Verification Task. The Multi-Modal Fact Verification task challenges MLLMs to verify the accuracy of claims using retrieved multi-modal evidence. In this task, the query $q$ can be a multi-modal claim, and the document collection $\mathcal{D}$ consists of both text and image documents, where the image documents do not contain captions. Each claim is assigned one of three labels, “Support”, “Refute”, or “Insufficient”, indicating whether the retrieved evidence supports, refutes or lacks sufficient information to verify the claim. We build this task on the Factify dataset (Mishra et al., 2022), but we focus on open-domain fact verification by retrieving evidence from a multi-modal document collection (Thorne et al., 2018).
@@
Image Reranking Task. In the Image Reranking task, the objective is to identify the most relevant images based on a given image description. In this setting, the image description serves as query $q$, and the document collection $\mathcal{D}$ consists of image documents without associated captions. For each description, we first use a multi-modal retriever to retrieve candidate image documents based solely on their image features and then rerank the images using MLLMs. To adapt MLLMs for this task, we follow previous work (Muennighoff, 2022) and compute the Perplexity (PPL) score to rerank image candidates based on their image features. This approach models the relevance between queries and images in a manner similar to image captioning, where a lower PPL score indicates greater relevance between the candidate image and the given query.
Table 1. Comparison of Multi-Modal Benchmarks.
Benchmarks Input Retrieval Multi Task Open Domain
MSCOCO (2014) Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Flickr30K (2014) Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
K-VQA (2019) Multi [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WebQA (2022) Text Multi [Uncaptioned image] [Uncaptioned image]
MRAG-Bench (2024) Multi Image [Uncaptioned image] [Uncaptioned image]
M2RAG (Ours) Multi Multi [Uncaptioned image] [Uncaptioned image]